OpenStudy (anonymous):

The wind-chill factor measures how cold it feels based on both temperature and wind speed. The wind-chill factor was 8 Farenheit at 6 p.m. and dropped to -8 farenheit by 10 p.m. What was the change in the wind-chill factor between 6 p.m. and 10 p.m.?

OpenStudy (mathstudent55):

You need to subtract 8 minus -8.

OpenStudy (anonymous):

thank you! i wanted to see if i was right so its 16 right?

OpenStudy (mathstudent55):

\(8 - (-8)\) \(= 8 + 8 \) \(= 16\)

OpenStudy (mathstudent55):

Yes, you are correct.

OpenStudy (anonymous):

thank you!

OpenStudy (mathstudent55):

You're welcome.